# Enhancing Accuracy in Differentially Private Distributed Optimization Through Sensitivity Reduction

### v1

@@ 6 Numerical Results
Figure 2: Comparison of convergence between Algorithm 1 and the DPOP method [12] under the same privacy budget.
@@
Figure 3: Comparison of convergence between Algorithm 1 and the PP-DOAGT method [16] under the same noise.
@@
6.1 Comparison with Other Differentially Private Methods
We consider a multi-agent system of 100 sensors, where the communication network is generated using an Erdős-Rényi random graph model with an edge creation probability of 0.1, as illustrated in Fig. 1. The dimensions in Problem (42) are set as $m=3$ and $p=2$. For Algorithm 1, the initial states $\bm{x}_{i}(0)$ are randomly sampled from a standard normal distribution. The stepsize $\alpha_{k}$ and noise parameter $\nu_{k}$ are set according to (5), and the relevant parameters $\gamma$, $\beta$, $q_{2}$, and $q_{1}$ are tuned following the heuristic strategy described in Remark 5.19. We implement Algorithm 1 under different privacy budgets ($\epsilon\in\{0.1,1,10\}$). For $\epsilon=0.1/1/10$, the parameters are set as $\gamma=0.001/0.001/0.002$, $\beta=1000/1000/100$, $q_{2}=0.99/0.99/0.99$, and $q_{1}=0.92/0.97/0.97$, respectively. For each case, we calculate the mean and standard deviation of the residuals (denoted as ${\|\mathbf{x}(k)-\mathbf{x}^{*}\|^{2}}$) over 100 trials, with the results shown as the red line in Fig. 2. It can be observed that our algorithm quickly approaches the optimal solution under all privacy budgets. This confirms that the proposed algorithm can achieve arbitrary $\epsilon$-DP. Moreover, higher privacy levels result in lower optimization accuracy, which aligns with the conclusion of Theorem 5.17.
For comparison, we implement the DPOP method from [12] under the same privacy budgets. The corresponding results are shown by the blue lines in Fig. 2. It can be seen that our algorithm achieves higher accuracy than the DPOP method. This is because Algorithm 1’s sensitivity is more compact, and the introduction of auxiliary variable $\bm{y}_{i}$ enhances its convergence. Additionally, we also compare the proposed algorithm with the latest PP-DOAGT method from [16] under the same noise conditions. The comparison results are shown in Fig. 3. At higher privacy levels (the privacy budgets of the PP-DOAGT method are 18446.44, 3652.67, and 365.27 with $\delta=1$, respectively), our algorithm achieves comparable optimization accuracy to the PP-DOAGT method. This is due to the fact that our algorithm has lower sensitivity than the PP-DOAGT method (as shown in Table 2).
6.2 Comparison with Other Types of Privacy-Preserving Methods
@@
[…]
@@ 7 Conclusion
In this paper, we propose a differentially private distributed optimization algorithm. Compared to existing state-of-the-art algorithms, the proposed algorithm exhibits lower sensitivity, thereby achieving higher optimization accuracy under the same privacy budget. Furthermore, the proposed algorithm reduces both the number of communication variables and the amount of injected noise by half, significantly enhancing communication and computational efficiency. Moreover, we rigorously analyze the algorithm’s differential privacy, establish its convergence, and derive an upper bound on its optimization accuracy. The theoretical results reveal the inherent trade-off between privacy and accuracy and provide guidance for parameter tuning. Future work will focus on developing differentially private distributed algorithms with lower sensitivity for directed graphs.
@@ 9 Proof of Basic Lemmas
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (50)
@@
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (51)
@@
$\quad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k+1)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (52)
@@
$\quad+\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
@@
$\quad-\alpha_{k}\nabla F(\mathbf{z}(k)).$
@@
$\quad\mathbb{E}\left[\langle\mathbf{x}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (54)
@@
$\quad\mathbb{E}\left[\langle\mathbf{1}\bar{\bm{x}}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (55)
@@
$\quad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$ (56)
@@
$\quad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]$ (57)
@@
$\quad+2\mathbb{E}\left[\langle\mathbf{x}(k),-\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$
@@
$\quad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (58)
@@
$\quad+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$
@@
$\quad+\frac{L^{2}\alpha_{k+1}^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2p\nu_{k+1}^{2}+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right].$
@@
$\quad 2\langle\bm{a}_{1},-\bm{b}_{1}\rangle$ (59)
@@
$\quad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2},$
@@
$\quad\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]$ (60)
@@
$\quad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right].$
@@
$\quad\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$ (61)
@@
$\quad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (62)
@@
$\quad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+(2p+4pL\alpha_{k+1})\nu_{k+1}^{2}$
@@
$\quad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2p(1+2L\alpha_{k+1}+\frac{L^{2}\alpha_{k+1}}{\mu})\nu_{k+1}^{2},$
@@
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (63)
@@
$\quad+\mathbb{E}\left[\|\bm{d}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$
$\quad+\mathbb{E}\left[2\langle\bm{b}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right].$
@@
$\quad\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}$ (66)
@@
$\quad+2\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\Big{\|}^{2}$
@@
$\quad\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]$ (67)
@@
$\quad+2n\alpha_{k+1}^{2}\mathbb{E}\Big{[}\Big{\|}\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big{\|}^{2}|\mathcal{M}_{k}\Big{]}$
@@
$\quad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
@@
$\quad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+12npL^{2}\alpha_{k+1}^{2}\nu_{k+1}^{2}+4npL^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}\nu_{k}^{2},$
@@
$\quad\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]$ (69)
@@
$\quad\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$ (70)
@@
$\quad+\frac{2\alpha_{k+1}}{\alpha_{k}}\mathbb{E}\left[\left\langle\mathbf{1}\bar{\bm{x}}(k),W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
@@
$\quad\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$ (74)
@@
$\qquad\Big{(}(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big{)}\bm{\xi}(k+1)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\Big{[}\Big{\langle}(\frac{1}{n}\mathbf{1}\mathbf{1}^{T}-I)\nabla F(\mathbf{z}(k+1)),-W\bm{\xi}(k)\Big{\rangle}|\mathcal{M}_{k}\Big{]}$
$\quad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),-W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
$\leq 12npL\alpha_{k+1}\nu_{k+1}^{2}+2npL(4+np+2L\alpha_{k})\alpha_{k+1}\nu_{k}^{2}.$
@@
$\quad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (75)
@@
$\quad+\frac{2+\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]+2np(9+6L\alpha_{k+1})\nu_{k+1}^{2}$
$\quad+2np(15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1})\nu_{k}^{2}$
@@
$\quad+\frac{4n(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\quad+\frac{(2+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\quad+2np\Big{(}9+6L\alpha_{k+1}+\frac{18L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big{)}\nu_{k+1}^{2}$
$\quad+2np\Big{(}15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$

### v2

@@ 6.2 Comparison of Differentially Private Methods Based on Distributed Learning on MNIST
Figure 6: Accuracy comparison between Algorithm 1 and DPOP [18] under the same privacy budget on MNIST image classification.
@@
Figure 7: Accuracy comparison between Algorithm 1 and PP-DOAGT [24] under the same noise on MNIST image classification.
Figure 8: Accuracy comparison between Algorithm 1 and Diff-TV-AB [21] under the same noise on MNIST image classification.
For Algorithm 1, the relevant parameters are set as $\gamma=0.001$, $\beta=100$, $q_{2}=0.999$, and $q_{1}=0.998$. We record the evolution of training and testing accuracy of Algorithm 1 under a privacy budget of $\epsilon=1$, as shown by the blue lines in Fig. 6. The results demonstrate that the final training and testing accuracy both exceed 90%, indicating that our algorithm can maintain high model performance while ensuring DP. For comparison, we implement the DPOP method from [18], with its performance depicted by the orange lines in Fig. 6. Under the same privacy budget, our algorithm demonstrates superior performance in both training and testing accuracy compared to the DPOP method. Furthermore, we reproduce two recently proposed methods, PP-DOAGT [24] and Diff-TV-AB [21], with experimental results shown by the orange lines in Figs. 7 and 8. The results indicate that under the same noise conditions, our algorithm achieves optimization accuracy comparable to both PP-DOAGT and Diff-TV-AB. Benefiting from its lower sensitivity design (as shown in Table 3), our algorithm exhibits stronger privacy protection performance relative to PP-DOAGT and Diff-TV-AB.
Figure 9: Testing accuracy of Algorithm 1 under different $\gamma$ and $\beta$ parameter configurations.
To evaluate the robustness and stability of Algorithm 1, we conduct a parameter sensitivity analysis, focusing on the effects of initial stepsize parameter $\gamma$ and parameter $\beta$ on model performance. The parameter $\gamma$ is tested over the range $\{0.001,0.005,0.01,0.05,0.1\}$, while $\beta$ is evaluated across $\{0.1,1,10,100,1000\}$. Experimental results using testing accuracy as the evaluation metric are presented in Fig. 9. Analysis reveals that when $\gamma$ takes relatively small values (specifically $\gamma\leq 0.01$), and $\beta$ satisfies $\gamma\beta\leq 1$, the algorithm achieves stable testing accuracy exceeding 85%, demonstrating desirable performance levels. This finding experimentally validates the necessity of the condition $\gamma\beta\leq 1$ for convergence. Furthermore, we observe that within the range of $\gamma\leq 0.01$, even with substantial variations in $\beta$ values, testing accuracy remains consistently high as long as $\gamma\beta\leq 1$ is maintained. This indicates Algorithm 1’s insensitivity to parameter variations and confirms its strong robustness.
Figure 10: Privacy-accuracy trade-off of Algorithm 1 under different network topologies and scales.
To further evaluate the performance of Algorithm 1, we systematically investigate its privacy-accuracy trade-off characteristics under different network topologies (ring graph and ER random graph) and network scales ($n=50$ and $n=100$). The experimental results are presented in Fig. 10. Analysis of Fig. 10 reveals that across all experimental configurations, testing accuracy consistently improves with increasing the privacy budget $\epsilon$, clearly demonstrating the inherent privacy-accuracy trade-off in differentially private algorithms, which aligns with the theoretical result in Theorem 5.18. Furthermore, under the same privacy budgets, the ER random graph topology achieves higher testing accuracy compared to the ring graph topology. This phenomenon can be attributed to the higher connectivity of random graphs, which facilitates information propagation among agents, thereby accelerating convergence and enhancing final performance. Notably, across all combinations of network topologies and scales, Algorithm 1 maintains testing accuracy above 85% when the privacy budget $\epsilon\geq 0.5$. These results demonstrate both the scalability of Algorithm 1 and its capability to maintain high model accuracy while providing moderate privacy protection.
@@
6.3 Comprehensive Performance Comparison Across Different Types of Privacy-Preserving Methods
@@
$\text{M-NMI}\triangleq\max_{1\leq k\leq K}\left\{\frac{I(\mathcal{V}_{i}(k),\mathcal{I}_{j}(k))}{I(\mathcal{V}_{i}(k),\mathcal{V}_{i}(k))}\right\},$ (43)
@@
Based on the comparative results presented in Table 4, the proposed algorithm demonstrates distinct advantages in the following practical deployment scenarios:
  1. 1.
    Scenarios Requiring Defense Against Both Internal and External Threats: Correlated randomness-based methods fail to protect against external eavesdroppers (M‑NMI = 1). The proposed algorithm achieves low M‑NMI values under the external eavesdropper and honest-but-curious adversary, confirming its capability to resist both types of adversaries simultaneously.
  2. 2.
    Scenarios Demanding Tunable and Quantifiable Privacy Guarantees: Both correlated randomness-based and encryption-based methods offer a fixed level of privacy protection. Our algorithm allows the privacy budget $\epsilon$ to be adjusted according to specific application requirements, providing configurable and provable privacy.
  3. 3.
    Resource-Constrained Large-Scale Distributed Systems: Encryption-based methods incur high computational and communication overhead due to frequent encryption and decryption operations. The proposed algorithm only requires lightweight noise injection and, compared to existing DP gradient-tracking methods, reduces the number of communication variables and noise injections by half, significantly lowering the resource burden.
  4. 4.
    Networks with Dynamic or Unknown Topology: Correlated randomness-based methods rely on the assumption that each honest agent has at least one honest neighbor, which is often impractical in self-organizing or time-varying networks. Our algorithm only requires network connectivity, making it directly applicable to mobile or peer-to-peer networks without imposing additional topological constraints.
In summary, the proposed differentially private algorithm provides a versatile and practical solution for privacy-preserving distributed optimization. It effectively balances provable privacy, a configurable privacy-accuracy trade-off, lower resource overhead, and broader topological applicability under various real-world constraints and threat models.
@@ 7 Conclusion
In this paper, we propose a differentially private distributed optimization algorithm for undirected graphs, and derive an explicit closed-form expression for the noise parameter as a function of the privacy budget. Compared to existing state-of-the-art algorithms, the proposed algorithm exhibits lower sensitivity, thereby achieving higher optimization accuracy under the same privacy budget. Furthermore, the proposed algorithm reduces both the number of communication variables and the amount of injected noise by half, significantly enhancing communication and computational efficiency. Moreover, we rigorously analyze the algorithm’s differential privacy, establish its convergence, and derive an upper bound on its optimization accuracy. The theoretical results reveal the inherent trade-off between privacy and accuracy and provide guidance for parameter tuning. Future work will proceed along three promising directions: (i) Extending the sensitivity‑reduction framework to directed graph topologies. (ii) Enhancing the algorithm’s convergence rate while preserving its privacy and accuracy guarantees, where potential pathways include designing adaptive stepsize schedules (e.g., two-phase or noise-aware decay) and incorporating variance‑reduction techniques. (iii) Generalizing the algorithm to broader problem classes, including convex (but not strongly convex) objectives and structured nonconvex problems frequently encountered in modern machine learning.
@@ 9 Proof of Basic Lemmas
$\qquad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (50)
@@
$\qquad\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (51)
@@
$\qquad\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k+1)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$ (52)
@@
$\qquad+\mathbb{E}\left[\left\langle\nabla F(\mathbf{x}(k)),\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
@@
$\qquad-\alpha_{k}\nabla F(\mathbf{z}(k)).$
@@
$\qquad\mathbb{E}\left[\langle\mathbf{x}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (54)
@@
$\qquad\mathbb{E}\left[\langle\mathbf{1}\bar{\bm{x}}(k),\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$ (55)
@@
$\qquad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$ (56)
@@
$\qquad\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{z}(k)\|^{2}|\mathcal{M}_{k}\right]$ (57)
@@
$\qquad+2\mathbb{E}\left[\langle\mathbf{x}(k),-\bm{\xi}(k)\rangle|\mathcal{M}_{k}\right]$
@@
$\qquad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (58)
@@
$\qquad+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$
@@
$\qquad+\frac{L^{2}\alpha_{k+1}^{2}}{n}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2p\nu_{k+1}^{2}+\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right].$
@@
$\qquad 2\langle\bm{a}_{1},-\bm{b}_{1}\rangle$ (59)
@@
$\qquad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2},$
@@
$\qquad\mathbb{E}\left[2\langle\bm{a}_{1},-\bm{b}_{1}\rangle|\mathcal{M}_{k}\right]$ (60)
@@
$\qquad+\frac{L^{2}\alpha_{k+1}(1-\alpha_{k+1}\mu)}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right].$
@@
$\qquad\mathbb{E}\left[2\langle-\bm{b}_{1},\bm{c}_{1}\rangle|\mathcal{M}_{k}\right]$ (61)
@@
$\qquad\mathbb{E}\left[\|\bar{\bm{x}}(k+1)-\bm{x}^{*}\|^{2}|\mathcal{M}_{k}\right]$ (62)
@@
$\qquad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+(2p+4pL\alpha_{k+1})\nu_{k+1}^{2}$
@@
$\qquad+\frac{L^{2}\alpha_{k+1}}{n\mu}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2p(1+2L\alpha_{k+1}+\frac{L^{2}\alpha_{k+1}}{\mu})\nu_{k+1}^{2},$
@@
$\qquad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (63)
@@
$\qquad+\mathbb{E}\left[\|\bm{d}_{2}\|^{2}|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$
$\qquad+\mathbb{E}\left[2\langle\bm{b}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle\bm{a}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$
$\qquad+\mathbb{E}\left[2\langle\bm{b}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]+\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right].$
@@
$\qquad\Big\|\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big\|^{2}$ (66)
@@
$\qquad+2\Big\|\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{1}\bar{\bm{x}}(k))\Big\|^{2}$
@@
$\qquad\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]$ (67)
@@
$\qquad+2n\alpha_{k+1}^{2}\mathbb{E}\Big[\Big\|\frac{1}{n}\bm{1}^{T}\nabla F(\mathbf{z}(k+1))\Big\|^{2}|\mathcal{M}_{k}\Big]$
@@
$\qquad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{z}(k+1)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
@@
$\qquad+4L^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\right\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+4nL^{2}\alpha_{k+1}^{2}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
[…]
@@
$\qquad\mathbb{E}\left[2\langle\bm{a}_{2},\bm{b}_{2}\rangle|\mathcal{M}_{k}\right]$ (69)
@@
$\qquad\mathbb{E}\left[2\langle\bm{a}_{2},-\bm{c}_{2}\rangle|\mathcal{M}_{k}\right]$ (70)
@@
$\qquad+\frac{2\alpha_{k+1}}{\alpha_{k}}\mathbb{E}\left[\left\langle\mathbf{1}\bar{\bm{x}}(k),W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
@@
$\qquad\mathbb{E}\left[2\langle-\bm{c}_{2},\bm{d}_{2}\rangle|\mathcal{M}_{k}\right]$ (74)
@@
$\hskip 18.49988pt\Big((1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\Big)\bm{\xi}(k+1)\Big\rangle|\mathcal{M}_{k}\Big]$
$\qquad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\Big[\Big\langle(\frac{1}{n}\mathbf{1}\mathbf{1}^{T}-I)\nabla F(\mathbf{z}(k+1)),-W\bm{\xi}(k)\Big\rangle|\mathcal{M}_{k}\Big]$
$\qquad+2\frac{\alpha_{k+1}^{2}}{\alpha_{k}}\mathbb{E}\left[\left\langle\nabla F(\mathbf{z}(k)),-W\bm{\xi}(k)\right\rangle|\mathcal{M}_{k}\right]$
[…]
@@
$\qquad\mathbb{E}\left[\|\mathbf{x}(k+1)-\mathbf{1}\bar{\bm{x}}(k+1)\|^{2}|\mathcal{M}_{k}\right]$ (75)
@@
$\qquad+\frac{2+\sigma^{2}}{1-\sigma^{2}}\mathbb{E}\left[\|\bm{c}_{2}\|^{2}|\mathcal{M}_{k}\right]+2np(9+6L\alpha_{k+1})\nu_{k+1}^{2}$
[…]
@@
$\qquad+\frac{4n(2+\sigma^{2})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\mathbb{E}\left[\left\|\bar{\bm{x}}(k)-\bm{x}^{*}\right\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+\frac{(2+\sigma^{2})(q_{1}^{2}+2L^{2}\alpha_{k+1}^{2})}{1-\sigma^{2}}\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}|\mathcal{M}_{k}\right]$
$\qquad+2np\Big(9+6L\alpha_{k+1}+\frac{18L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big)\nu_{k+1}^{2}$
[…]
$\hskip 18.49988pt+\frac{6L^{2}(7+2L\alpha_{k})\alpha_{k+1}^{2}}{1-\sigma^{2}}\Big)\nu_{k}^{2}.$